A particle at rest is dropped under gravity from height h and it travels a distance 9h/25 in the last second find the height
 

Grade:11

1 Answers

Arun
5 years ago
Dear Prabhat
 
As the particle is dropped then u = 0
        As we all know that g = 9.8 m/s^2.
Then from the third equation of motion=
S = ut+1/2 gt^2
S = ½ gt^2
where t is the total time taken
  • Distance covered in (t-1) sec = h -9h/25 = 16h/25
Solving (1.) and (2.)
T = 5s
Substituting the value of t in eq. 1
S = ½ *9.8 *25
S = 122.5 m